OpenStudy (anonymous):

Quincy had a coupon for his favorite clothing store. He received a discount of sixteen dollars on a pair of jeans that originally cost forty-eight dollars. Set up a percent proportion to determine the percent discount he received with his coupon. Which of the following is a numerator in the proportion?

OpenStudy (anonymous):

sixteen thirty-two forty-eight sixty-four

OpenStudy (anonymous):

ok so 16 dollar discount on 48 dollar jeans

OpenStudy (anonymous):

16/48=3

OpenStudy (anonymous):

so he got 33% off

OpenStudy (anonymous):

or 1/3 of the price off

OpenStudy (anonymous):

x/100 = 16/48 cross multiply 48x = 1600 x = 1600/48 x = 33.3%

OpenStudy (anonymous):

which is not an answer choice....

OpenStudy (anonymous):

oh didnt see the rest of the question

OpenStudy (anonymous):

waht do they mean by proportion

OpenStudy (anonymous):

of i c, kellie got it right 48

OpenStudy (anonymous):

the question asker already left lol

OpenStudy (anonymous):

It is asking for the numerator, not the denominator

OpenStudy (anonymous):

it should be 33, but that is not an answer choice